OpenStudy (anonymous):

Do I have the answer? Six 6-sided dice are rolled. What is the probability that the number of dice showing even numbers and the number of dice showing odd numbers is equal?

OpenStudy (anonymous):

Can someone please confirm? I just used the binomial theorem

OpenStudy (anonymous):

@satellite73 Can I even use the binomial theorem here?

OpenStudy (anonymous):

I think it would be 20*(1/2)^6?

OpenStudy (anonymous):

In other words, 6c3*(1/2)^3*(1/2)^3

OpenStudy (mathmate):

Yes, this is a binomial distribution. So P(3)=C(6,3)(1/2)^3(1/2)^3 =20(1/2)^6 as you had it.

OpenStudy (anonymous):

Its correct, thanks.

OpenStudy (mathmate):

You're welcome!